# DarkDNS: Revisiting the Value of Rapid Zone Update

### v2

@@ Abstract.
Malicious actors exploit the DNS namespace to launch spam campaigns, phishing attacks, malware, and other harmful activities. Combating these threats requires visibility into domain existence, ownership and nameservice activity that the DNS protocol does not itself provide. To facilitate visibility and security-related study of the expanding gTLD namespace, ICANN introduced the Centralized Zone Data Service (CZDS) that shares daily zone file snapshots of new gTLD zones. However, a remarkably high concentration of malicious activity is associated with domains that do not live long enough make it into these daily snapshots. Using public and private sources of newly observed domains, we discover that even with the best available data there is a considerable visibility gap in detecting short-lived domains. We find that the daily snapshots miss at least 1% of newly registered and short-lived domains, which are frequently registered with likely malicious intent. In reducing this critical visibility gap using public sources of data, we demonstrate how more timely access to TLD zone changes can provide valuable data to better prevent abuse. We hope that this work sparks a discussion in the community on how to effectively and safely revive the concept of sharing Rapid Zone Updates for security research. Finally, as a contribution of this work, we are releasing a public live feed of newly registered domains, with the aim of enabling further research in early abuse identification.
@@ 1. Introduction
This inherent opacity has motivated many attempts to improve visibility into the DNS ecosystem to facilitate research, analysis, and improved operational security. A prominent example is the ICANN Centralized Zone Data Service (CZDS) which coordinates collection and restricted sharing of daily zone file snapshots of participating TLDs. However, previous studies have found that the 24-hour zone files, while an unprecedented enabler of DNS-related security research, leave vital visibility gaps in studies of abuse and exploits (Akiwate et al., 2020; Akiwate et al., 2021; Security and Stability Advisory Committee, 2024). Since so much malicious activity often begins soon after a domain is registered for that purpose, and the CZDS snapshots are updated only daily, commercial threat intelligence services (Domain Tools, 2024) have tried to fill the gap by detecting domains soon after registration.
@@ 2. Related Work
Barron et al.  (Barron et al., 2019) found significant correlation between early domain name deletions and potentially malicious activities, and noted that often zone files do not reveal whether a domain is registered. Other studies have found that spammer domains generally had short lifetimes, and that blocklists generally did not detect them until they were actively used and reported rather than when initially registered (Hao et al., 2013). Affinito et al.  (Antonia Affinito and Raffaele Sommese and Gautam Akiwate and Stefan Savage and Kimberly Claffy and Geoffrey M. Voelker and Alessio Botta and Mattijs Jonker, 2022) found patterns of suspiciously short lifetimes domain names in some TLDs (e.g., 80% under .xyz), many of which were reported as malicious. Their study revealed that domains registered under some TLDs (e.g., .xyz, .icu) are taken down within a few days, while others (e.g., .top) remain in the zone longer before removal.
@@
In the spike of COVID-related fraudulent online activity, a group of threat intelligence providers temporarily volunteered to share threat data in hopes of achieving better coverage. They found that for novel abuse types, of which there were many during COVID, the aggregated threat intelligence detected signals that registrars and registries did not (Bouwman et al., 2022).
A recent study is relevant to our goal of detection of newly observed domains without access to proprietary data. Last year, Sommese et al.  (Sommese et al., 2024). demonstrated that CT transparency logs are a reliable and up-to-date source of domain names for addressing the visibility gaps in ccTLDs, most of which do not participate in CZDS. They used CT logs to independently reconstruct more than half of the domains in several ccTLDs for which they had the full zone file as ground truth. Their work focused on spatial coverage whereas we use CT logs to analyze temporal gaps, i.e., how quickly we detect new and transient domains (Sommese et al., 2024).
@@ 4. Results
When comparing our list of detected newly registered domains to the diff between two daily CZDS zone snapshots, we found that our methodology identified 42% of the newly registered domains before they were published in the zone snapshots in our 3-month observation period. The remaining 58% did not have certificates issued immediately. Since our goal is to identify transient domains i.e., , domains that do not show up in zone snapshots, this comparison is purely functional.
@@ 4.2. Transient Domains
While we cannot investigate causes (i) and (ii) due to lack of RDAP data, we examined several domains in the third category, contacting the CERT teams of the CAs that issued certificates for these non-existent domains. GlobalSign, Sectigo, and Cloudflare confirmed that issuing a certificate for a non-existent domain (more precisely, issuing a certificate without validating the existence of the domain) is allowed under the condition that the CA possesses a previously obtained and valid Domain Validation (DV) token (see §3 and footnote 2). A necessary but not sufficient condition for this case is that the domain existed in the past. We conducted a comparison of the 34% of transient domains for which we failed to collect RDAP data with the CAIDA DZDB historical zone collection (Foster and Koga, 2024) and found that approximately 97% of those domains were registered in the past. Another possible approach to filter out these domains is to individually check if a certificate was issued in the DV token validity period in the past. We leave the complexity of longitudinal investigation of CT Logs as future work.
@@ Acknowledgments
We thank our anonymous shepherd and IMC reviewers for their insightful suggestions and feedback. We thank Cristian Hesselman and Roland van Rijswijk-Deij for their support and insights over time. Additionally, we extend our thanks to Domain Tools and SIDN Labs for providing us with access to passive DNS feeds and registry ground truth. This material is based on research sponsored by the National Science Foundation (NSF) grant OAC-2131987. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of NSF. This research was made possible by OpenINTEL, a joint project of the University of Twente, SURF, SIDN, and NLnet Labs; and operational support from the Twente University Centre for Cybersecurity Research (TUCCR).

### v4

@@ Abstract.
Malicious actors exploit the DNS namespace to launch spam campaigns, phishing attacks, malware, and other harmful activities. Combating these threats requires visibility into domain existence, ownership and nameservice activity that the DNS protocol does not itself provide. To facilitate visibility and security-related study of the expanding gTLD namespace, ICANN introduced the Centralized Zone Data Service (CZDS) that shares daily zone file snapshots of new gTLD zones. However, a remarkably high concentration of malicious activity is associated with domains that do not live long enough make it into these daily snapshots. Using public and private sources of newly observed domains, we discover that even with the best available data there is a considerable visibility gap in detecting short-lived domains. We find that the daily snapshots miss at least 1% of newly registered and short-lived domains, which are frequently registered with likely malicious intent. In reducing this critical visibility gap using public sources of data, we demonstrate how more timely access to TLD zone changes can provide valuable data to better prevent abuse. We hope that this work sparks a discussion in the community on how to effectively and safely revive the concept of sharing Rapid Zone Updates for security research. Finally, we release a public live feed of newly registered domains, with the aim of enabling further research in abuse identification.
DNS, RZU, transparency
@@ 1. Introduction
This inherent opacity has motivated many attempts to improve visibility into the DNS ecosystem to facilitate research, analysis, and operational security. A prominent example is the ICANN Centralized Zone Data Service (CZDS) which coordinates collection and restricted sharing of daily zone file snapshots of participating TLDs. However, previous studies have found that the 24-hour zone files, while an unprecedented enabler of DNS-related security research, leave vital visibility gaps in studies of abuse and exploits (Akiwate et al., 2020, 2021; Security and Stability Advisory Committee, 2024). Since so much malicious activity often begins soon after a domain is registered for that purpose, and the CZDS snapshots are updated only daily, commercial threat intelligence services (Domain Tools, 2024) have tried to fill the gap by detecting domains soon after registration.
@@ 2. Related Work
Barron et al.  (Barron et al., 2019) found significant correlation between early domain name deletions and potentially malicious activities. Other studies found that spammer domains generally had short lifetimes, and that blocklists generally did not detect them until they were actively used and reported rather than when initially registered (Hao et al., 2013). Affinito et al.  (Antonia Affinito and Raffaele Sommese and Gautam Akiwate and Stefan Savage and Kimberly Claffy and Geoffrey M. Voelker and Alessio Botta and Mattijs Jonker, 2022) found patterns of suspiciously short-lived domain names in some TLDs (e.g., 80% under .xyz), many of which were reported as malicious. Their study revealed that domains registered under some TLDs (e.g., .xyz, .icu) are taken down within a few days, while others (e.g., .top) remain in the zone longer.
@@
In the spike of COVID-related fraudulent online activity, a group of threat intelligence providers temporarily volunteered to share threat data in hopes of achieving better coverage. They found that for novel abuse types, the aggregated threat intelligence detected signals that registrars and registries did not (Bouwman et al., 2022).
A recent study of Sommese et al.  (Sommese et al., 2024) is relevant to our goal of detection of newly observed domains without access to proprietary data. They demonstrated that CT logs are a reliable and up-to-date source of domain names for addressing the visibility gaps in ccTLDs, most of which do not participate in CZDS. They used CT logs to independently reconstruct more than half of the domains in several ccTLDs for which they had the full zone file as ground truth. Their work focused on spatial coverage whereas we use CT logs to analyze temporal gaps, i.e., how quickly we detect new and transient domains.
@@ 4. Results
When comparing our list of detected newly registered domains to the diff between two daily CZDS zone snapshots, we found that our methodology identified 42% of the newly registered domains before they were published in the zone snapshots in our 3-month observation period. The remaining 58% did not have certificates issued immediately. Since our goal is to identify transient domains i.e., domains that do not show up in zone snapshots, this comparison is purely functional.
@@ 4.2. Transient Domains
While we cannot investigate causes (i) and (ii) due to lack of RDAP data, we examined several domains in the third category, contacting the CERT teams of the CAs that issued certificates for these non-existent domains. GlobalSign, Sectigo, and Cloudflare confirmed that issuing a certificate for a non-existent domain (more precisely, issuing a certificate without validating the existence of the domain) is allowed under the condition that the CA possesses a previously obtained and valid Domain Validation (DV) token (see §3 and footnote 2). A necessary but not sufficient condition for this case is that the domain existed in the past. We conducted a comparison of the 34% of transient domains for which we failed to collect RDAP data with the CAIDA DZDB historical zone collection (Foster and Koga, 2024) and found that approximately 97% of those domains were registered in the past. Another possible approach to filter out these domains is to individually check if a certificate was issued in the DV token validity period in the past. Given the complexity of longitudinal investigation of CT Logs, we leave this filtering as future work.
@@ Acknowledgments
We thank our anonymous shepherd and IMC reviewers for their insightful suggestions and feedback. We thank Cristian Hesselman and Roland van Rijswijk-Deij for their support and insights over time. Additionally, we extend our thanks to Domain Tools and SIDN Labs for providing us with access to passive DNS feeds and registry ground truth. This material is based on research sponsored by the National Science Foundation (NSF) grant OAC-2131987. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of NSF. This work was partially supported by the GÉANT GN5-1 programme funded by the European Commission. This research was made possible by OpenINTEL (van Rijswijk-Deij et al., 2016), a joint project of the University of Twente, SURF, SIDN, and NLnet Labs; and operational support from the Twente University Centre for Cybersecurity Research (TUCCR).